|
ID |
Date |
Author |
Subject |
|
571
|
Thu May 18 02:16:05 2017 |
CB, TD, OH | Thursday 18 May | 10.10 DAQ continues OK
Current file May2017/R10_346
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
10.20 E17/F11 ambient temperature +24.3 deg C, RH 39.0%, d.p. +9.3 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
10.25 DAQ stops OK
Current file May2017/R13_294
Files May2017/R13* *deleted*
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
BNC PB-4 settings
Amplitude 90,000
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate 250Hz range
delay MIN, tail pulse
Check ASIC Control x2
Zero all spectra
10.30 DAQ starts
File May2017/R14
17.34 DAQ stops OK
Current file May2017/R14_113
Files May2017/R14* *deleted*
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
BNC PB-4 settings
Amplitude 90,000
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate 250Hz range
delay MIN, tail pulse
Check ASIC Control
Zero all spectra
*all* system wide checks OK
17.37 DAQ starts
File May2017/R15
Detector biases & leakage currents OK - see attachments 6 & 7
FEE64 Temperatures OK - see attachment 8
good events statistics - higher rates for several FEE64s - due to addition of system consoles? - see
attachment 9
No warnings/errors reported by merger since DAQ GO - ~ +30% data rate - see attachment 10 |
|
570
|
Wed May 17 05:28:19 2017 |
TD, OH | Wednesday 17 May | 11.11 DAQ found stopped
Most recent file May2017/R11_245 (last accessed 06:09)
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics zero - see attachment 4
Merger and Tapeserver terminal server windows are gone - see attachment 5
*all* system wide checks OK
No relevant system log messages (/var/log/messages) c. 06.00
DAQ STOP attempt #1 did not stop
#2 timeout nnaida1, nnaida2-24 did not stop
#3 timeout nnaida1, nnaida2-24 did not stop
Restart, configure and setup/go merger terminal server
Restart TapeServer terminal server
DAQ STOP attempt #3 all stop OK
11.24 Analysis of data file R11_245 - see attachment 6
dead-time/idle-time < 0.5%
AIDA SYNC100 variations ~160/78000
11.34 E17/F11 ambient temperature +24.5 deg C, RH 37.0%, d.p. +9.0 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
13.06 Check ASIC Control x2
Zero all spectra
Restart terminal servers for Httpd for Merger/TapeServer, Merger & TapeServer
Restart browser for Merger & TapeServer
N.B. *no FEE64 power cycle, no DAQ RESET, SETUP etc*
11.30 DAQ starts
File May2017/R12
13.10 DAQ GO - data transfer links for nnaida11 & nnaida12 dropped
DAQ STOP
Disable all data transfer links
Restart terminal servers for Httpd for Merger/TapeServer, Merger & TapeServer
Restart browser for Merger & TapeServer
Enable all data transfer links
DAQ GO OK
File May2017/R13
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff |
|
568
|
Tue May 16 03:18:46 2017 |
TD, OH | Tuesday 16 May | 11.19 DAQ continues OK
Current file May2017/R10_346
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
11.22 Analysis of data file R10_246 - see attachment 6
dead-time/idle-time < 1%
AIDA SYNC100 variations ~700/97000
10.10 E17/F11 ambient temperature +24.5 deg C, RH 38.0%, d.p. +9.4 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
11.23 DAQ stops OK
Current file May2017/R10_347
Files May2017/R10* *deleted*
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
BNC PB-4 settings
Amplitude 90,000
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate 250Hz range
delay MIN, tail pulse
Check ASIC Control x2
Zero all spectra
11.30 DAQ starts
File May2017/R11
18.18 Detector biases & leakage currents OK - see attachments 7 & 8
FEE64 Temperatures OK - see attachment 9
18.24 E17/F11 ambient temperature +24.5 deg C, RH 36.8%, d.p. +8.7 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%) |
|
567
|
Mon May 15 02:01:02 2017 |
TD, OH | Monday 15 May | 10.01 DAQ continues OK
Current file May2017/R8_334
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
10.06 Analysis of data file R8_333 - see attachment 6
dead-time/idle-time < 2%
AIDA SYNC100 variations ~1800/95000
10.10 E17/F11 ambient temperature +24.8 deg C, RH 36.4%, d.p. +8.9 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
10.15 DAQ stops OK
Current file May2017/R7_298
Files May2017/R8* *deleted*
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator *0x08*
LEC/MEC fast comparator 0xff
BNC PB-4 settings
Amplitude 90,000
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate 250Hz range
delay MIN, tail pulse
Check ASIC Control x2
Zero all spectra
10.20 DAQ starts
File May2017/R8
Transient timeouts at DAQ GO for FEE64s nnaida1 and nnaida2 which resolved themselves without user
intervention
DAQ run control transitioned from 'RESET and STOP/GO' to 'RESET' to 'RESET and STOP/GO'
10.22 DAQ stops
File May2017/R9
Tapeserver data rates fluctuate 10-30Mb/s possibly due to parallel disk I/O activity (deleting R8 files)
Transient timeouts at DAQ STOP for FEE64s nnaida1 and nnaida2 which resolved themselves without user
intervention - see attachment 7
DAQ run control transitioned from 'RESET and STOP/GO' to 'RESET' to 'RESET and STOP/GO'
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator *0x0a*
LEC/MEC fast comparator 0xff
10.31 DAQ starts
File May2017/R10 |
|
566
|
Sun May 14 01:52:49 2017 |
TD, OH | Sunday 14 May | 09.54 DAQ continues OK
Current file May2017/R7_292
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
10.02 Rate/HitRate spectra - see attachments 6 & 7
1.8.L ADC spectra - see attachments 8 & 9
1.8.W waveforms - see attachments 10-13
10.07 Analysis of data file R7_294 - see attachment 14
dead-time/idle-time < 0.5%
AIDA SYNC100 variations ~330/102000
10.17 E17/F11 ambient temperature +25.1 deg C, RH 37.2%, d.p. +9.4 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
10.19 DAQ stops OK
Current file May2017/R7_298
Files May2017/R7* *deleted*
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
BNC PB-4 settings
Amplitude 90,000
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate 250Hz range
delay MIN, tail pulse
Check ASIC Control x2
Zero all spectra
10.22 DAQ starts
File May2017/R8
Analysis of May2017/R8_3 confirms ADC data synchronisation - see attachments 15 & 16 |
|
565
|
Sat May 13 03:16:16 2017 |
TD | Analysis of May2017/R1_0 to May2017/ R1_79 |
See https://elog.ph.ed.ac.uk/AIDA/555
Date files May2017/R1_0 to May2017/R_79 are an extended (c. 3.8 day) background run with
BNC PB-4 Pulser OFF and and a relatively high slow comparator threshold 0x64 to reduce
data volume.
Attachment 1 - DSSSD#1-6 m_p versus m_n
Attachment 2 - DSSSD#1-6 x versus y hit patterns
Attachment 3 - DSSSD#1-6 E_p versus E_n (dispersion ~20-23keV/channel)
Attachment 4 - DSSSD#1-6 E_p versus E_n (with tags)
Attachment 5 - 2D contour levels & colours for all 2D spectra above
Attachment 6 - MIDASsort subroutine
Attachment 7 - MIDASsort namelist I/O variables
Attachments 8-10 - ADC offset calculation logs etc
Attachment 11 - spectrum titles
MIDASsort calculates channel and ADC data as follows
ch = channel + ( module - 1 ) * 64 + ( range * 2048 )
adc_data = INT( RSHIFT( ABS( adc_data - 32768 ), 3 ) - offset(ch) + 0.5 )
where
channel = FEE channel # (0-63)
module = FEE module # (1-24)
range = LEC or HEC (LEC = 0, HEC = 1)
ADC offsets were calculated using the pulser walkthrough data of files May2017/R4
and command
cd ~/td/calibration
offsets /homes/npg/Spectra/R4 100 1400 15 10
where the peakfind peak search parameters are limits 100-1400 channels, fwhm=15 channels and acceptance=10%
Probably ~10% of ADC offsets need manual checking/adjustment
The 2D spectra (attachments 1-4) were produced for events with
multiplicities 0 < m_p < 8 *and* 0 < m_n < 8
all x-y combinations incremented
no clustering, gain matching or software thresholds
All DSSSDs show Uranium 4n+2 decay sequence background spectra from the intrinsic activity of the DSSSDs
and airborne decay products (e.g. 222Rn). Not clear why statistics are higher for DSSSD#1 cf. DSSSD#2-6.
One possibility is the activity of the aluminized Mylar window just upstream of DSSSD#1? Surrounding air
and Al snout would presumably affect DSSSD#1 and #6 about the same. |
|
564
|
Sat May 13 02:35:53 2017 |
TD, OH | Saturday 13 May | 10.35 DAQ continues OK
Current file May2017/R6_246
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x20
LEC/MEC fast comparator 0xff
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
10.46 Rate/HitRate spectra - see attachments 6 & 7
1.8.L ADC spectra - see attachments 8 & 9
1.8.W waveforms - see attachments 10-13
10.51 Analysis of data file R6_245 - see attachment 14
Zero AIDA pause/resume
Zero idle-time
AIDA SYNC100 variations ~33/503400
12.03 E17/F11 ambient temperature +24.6 deg C, RH 39.4%, d.p. +9.9 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
12.04 DAQ stops OK
Current file May2017/R6_250
Files May2017/R6* *deleted*
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x0a
LEC/MEC fast comparator 0xff
BNC PB-4 settings
Amplitude 90,000
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate 250Hz range
delay MIN, tail pulse
Check ASIC Control x2
Zero all spectra
12.15 DAQ starts
File May2017/R7
14.30 Analysis of May2017/R7_29
ADC data synchronised - see attachments 15 & 16
14.37 Analysis of data file R7_29 - see attachment 17
AIDA pause/resume for some FEE64s
dead-time/idle-time <0.5%
AIDA SYNC100 variations ~300/100000 |
|
563
|
Fri May 12 02:55:10 2017 |
TD, OH | Friday 12 May | 10.55 DAQ continues OK
Current file May2017/R6_181
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x20
LEC/MEC fast comparator 0xff
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
All histograms zero'd
11.04 Analysis of data file R6_180 - see attachment 6
Zero AIDA pause/resume
Zero idle-time
AIDA SYNC100 variations ~22/499760
11.20 E17/F11 ambient temperature +24.6 deg C, RH 39.6%, d.p. +9.9 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%) |
|
562
|
Thu May 11 05:43:30 2017 |
TD, OH | Thursday 11 May | 13.43 DAQ continues OK
Current file May2017/R6_123
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x20
LEC/MEC fast comparator 0xff
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
13.37 E17/F11 ambient temperature +24.4 deg C, RH 37.3%, d.p. +9.1 deg C
13.58 1.8.L ADC spectra - see attachments 6 & 7
1.8.W waveforms - see attachments 8-11
14.16 Analysis of data file R6_124 - see attachment 12
Zero AIDA pause/resume
Zero idle-time
AIDA SYNC100 variations ~20/468430
18.36 DAQ continues OK
Current file May2017/R6_137
Detector biases & leakage currents OK - see attachments 13 & 14
FEE64 Temperatures OK - see attachment 15
'good events' statistics OK - see attachment 16
*all* system wide checks OK |
|
561
|
Wed May 10 02:55:13 2017 |
TD, OH | Wednesday 10 May | 10.55 DAQ continues OK
Current file May2017/R6_44
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x20
LEC/MEC fast comparator 0xff
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
10.48 E17/F11 ambient temperature +24.7 deg C, RH 35.3%, d.p. +8.3 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
18.32 DAQ continues OK
Current file May2017/R6_67
Detector biases & leakage currents OK - see attachments 6 & 7
FEE64 Temperatures OK - see attachment 8
'good events' statistics OK - see attachment 9
*all* system wide checks OK |
|
560
|
Tue May 9 02:25:04 2017 |
TD, OH | Tuesday 9 May | 10.25 DAQ continues OK
Current file May2017/R1_79
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x64
LEC/MEC fast comparator 0xff
BNC PB-4 pulser OFF
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
10.28 1.8.L ADC spectra - see attachments 6 & 7
1.8.W waveforms - see attachments 8-11
10.43 DAQ stopped
Current file May2017/R1_79
10.48 E17/F11 ambient temperature +24.5 deg C, RH 29.0%, d.p. +5.0 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
10.58 DAQ starts
File May2017/R2 *IGNORE*
Pulser walkthough
BNC PB-4 settings
Amplitude 90,000 - 10,000 @ 10,000 step
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate ~250Hz
delay MIN, tail pulse
ASIC settings 2017Mar27-16.12.39
Slow comparator 0x0a
See attachment 12
11.02 DAQ starts
File May2017/R3 *IGNORE*
Pulser walkthough
BNC PB-4 settings
Amplitude 12,000 - 2,000 @ 2,000 step
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate ~250Hz
delay MIN, tail pulse
ASIC settings 2017Mar27-16.12.39
See attachment 13
Pulser walkthroughs R2 & R3 should be ignored as + polarity test input
had significantly lower amplitude (x2 oor more) than - polarity test input.
The + polarity test input is derived from tee'd output from BNC PB-4 and
EG&G Ortec 433 Sum & Invert - physically moving the cabling and connectors
between these two points changed tail pulse amplitude or produced a transient
(non-tail) pulse with duration ~ 100ns. Not clear which cable or connector
is the issue (all cables and connectors were individually replaced) but output
amplitude currently correct so will proceed.
11.22 DAQ starts
File May2017/R4
Pulser walkthough
BNC PB-4 settings
Amplitude 90,000 - 10,000 @ 10,000 step
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate ~250Hz
delay MIN, tail pulse
ASIC settings 2017Mar27-16.12.39
See attachment 14
11.28 DAQ starts
File May2017/R5
Pulser walkthough
BNC PB-4 settings
Amplitude 12,000 - 2,000 @ 2,000 step
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
rate ~250Hz
delay MIN, tail pulse
ASIC settings 2017Mar27-16.12.39
See attachment 15
17.25 E17/F11 ambient temperature +24.7 deg C, RH 28.1%, d.p. +5.0 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%)
17.32 Detector biases & leakage currents OK - see attachments 16-17
FEE64 Temperatures OK - see attachment 18
17.40 DAQ starts
File May2017/R6
Continuing DAQ test
BNC PB-4 settings
Amplitude 90,000
x5 attenuator IN
t_r 50ns, tau_d 50us
polarity -
frequency 25Hz range
delay MIN, tail pulse
ASIC settings 2017Mar27-16.12.39
slow comparator 0x20 |
|
559
|
Mon May 8 10:00:02 2017 |
Patrick Coleman-Smith | [HowTo] Operating document for the Pi_Monitor FEE64 console logging equipment | The operating proceedure for the Raspberry Pi FEE64 console monitor is attached.
Please let me know if more detail is required.
|
|
558
|
Mon May 8 02:11:37 2017 |
OH, TD | Monday 8 May | 10.11 DAQ continues OK
Current file May2017/R1_58
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x64
LEC/MEC fast comparator 0xff
BNC PB-4 pulser OFF
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
10.19 1.8.L ADC spectra - see attachments 6 & 7
1.8.W waveforms - see attachments 8-11
10.35 E17/F11 ambient temperature +24.2 deg C, RH 33.2%, d.p. +7.0 deg C
Julabo FL11006 set point +20 deg C, actual +20.0 deg C, water level OK (c. 65%) |
|
557
|
Sun May 7 01:29:50 2017 |
OH, TD | Sunday 7 May |
09.28 DAQ continues OK
Current file May2017/R1_36
ASIC settings 2017Mar27-16.12.39 *except*
slow comparator 0x64
LEC/MEC fast comparator 0xff
BNC PB-4 pulser OFF
Detector biases & leakage currents OK - see attachments 1 & 2
FEE64 Temperatures OK - see attachment 3
'good events' statistics OK - see attachment 4
No warnings/errors reported by merger since DAQ GO - see attachment 5
*all* system wide checks OK
09.40 1.8.L ADC spectra - see attachments 6 & 7
1.8.W waveforms - see attachments 8-11
10.00 E17/F11 ambient temperature +24.3 deg C, RH 39.4%, d.p. +9.6 deg C
18:11 System wide checks all ok
Leakage and bias all ok - see attachment 12
Fee temperatures all ok - see attachment 13
LN2 Refill cycle has been running the past few days and we have encountered no issues
AIDA daq continues to run fine
19:20 Ambient temperature 24.5 Humdity 38.3% D.P +9.2 |
|
556
|
Sat May 6 03:32:38 2017 |
OH, TD | Saturday 6 May |
11.30 Detector biases & leakage currents OK - see attachments 1 & 2
Ambient Temperature 24.3 Humidity 37.5% Dew Point 8.9
FEE64 temperatures - see attachment 3
Good events statistics (N.B. slow comparator=0x64, fast comparator=0xff) - see attachment 4
System wide checks all ok
Screen shots for layout 3 and 4 showing 1.8.L ADC spectra. Note that in the good events statistics a
higher rate is seen in the p+ junction strips
This is seen in the layouts as well.
16:00 Biases and leakages ok (attachment 5)
Ambient Temperature 24.4 Humidity 39.7 Dew Point 9.7
FEE64 Temperatures - attachment 6 |
|
555
|
Fri May 5 08:03:36 2017 |
OH, TD | Friday May 5th | Relatively painless start-up. Had an issue with firefox where all bookmarks had been forgotten.
All drop down menues in the DAQ did not work.
Kill all firefox processes to restore.
Had to restart DHCP to get the rasp pi working again.
At first no pulser was visible in all p-side FEEs. Cable and BNC adapter on output of inverter were switched out
and the inverter output
was switched from B to A.
After this pulser signals were visible in all ASICS bar NNAIDA 14 ASIC 1. Power cycle was needed to regain this.
No ADC data and no fast discriminator data. Waveform showed baseline at 16K compared to 8K expected.
Layouts 1,2,3,4,7 and 8 attached (Zoomed in waveforms are included for 7 and 8 around baseline) attachments 1-8
Temperatures included as attachment 9
Leakages and bias all ok see attachment 10
Ambient T 24.3C Humidity 34.5% Dew point 7.6C (Measurements taken AM)
system wide checks all ok.
Pulser settings
Amplitude 90000
x5 Attenuator in
Rise time tr 50ns fall time tau d 50us
Negative polarity
freq 25Hz range
Delay min
Internal ref, tail pulse
ASIC settings DB key 2017Mar27-16.12.39
16:26 Slow comparator threshold moved to 100
Fast comparator theshold LEC/MEC to 255
Run started to disk May2017/R1 |
|
554
|
Thu Apr 13 15:08:31 2017 |
CB | Card test | I tested the FWHM of a BNC PB-5 pulser signal as a function of shaping time using the first produced ISOL FEE64 card in Daresbury Labs.
Find attached four representative spectra per ASIC taken with 8, 5 and 0.5 us shaping time values. Run times are different.
Also find attached a plot of the FWHM as a function of the shaping time for the first channel in each ASIC. There appears to be a maximum of the FWHM at around 4 us. This may indicate the main source of noise is external, and at ~4 us the band pass lets most of the noise in.

Unlike the FEE cards employed in RIKEN, this card is grounded to the metal base plate through the screws. In order to investigate the effect of this ground on the noise level of the ADCs in the card, the card was disconnected from the metal plate base. An additional layer of insuating material was placed between the base and the FEE64 card. Nylon screws rather than metal screws were employed to affix the card to the base, and the same tests as before were repeated. See plot below. Overall, the same trend is recovered but suprisingly the FWHM is larger than before for ASICs #1,2 and #3. ASIC #4 has a lower FWHM than before. The reason for the change is not obvious, and may need further investigation.
Carlo

BNC PB-5 parameters
Delay 250 ns
Rate 100 Hz
Polarity POS
Pulse Top TAIL
Amplitude 0.5 V
|
|
553
|
Wed Apr 5 06:11:08 2017 |
OH | BigRIPS particle ID | I've been going through collecting the BigRIPS root files with Phong and making sure they are on AIDAs1.
The files we have for NP1306 are not the standard format so to get particle ID you need to do this command taken
from the RIKEN elog.
Command to construct the PID in the merged files from preliminary data:
Example file a11111_b1181_e_1130_ion.root
tree->Draw("(F11ICADC[0]+F11ICADC[1]+F11ICADC[2]+F11ICADC[3]+F11ICADC[4]+F11ICADC[5])/6:F11Pla_TM-F3Pla_TM+F5X*0.0383261>>h(500,0,25,500,0,16000)","","colz");
Though these root files were not created by the BigRIPS team so can be cleaned up further but should be good
enough for initial analysis. |
|
Draft
|
Wed Apr 5 06:10:55 2017 |
OH | BigRIPS particle ID | |
|
551
|
Sun Apr 2 04:40:00 2017 |
DK | BigRIPS Summary Info | It is all saved locally file:///homes/npg/Documents/2017Ca/
Although, a little more hacking is required to get the view.html webpage to allow you to click each run
Whether or not it is useful I don't know, but it was generated by a shell script
/homes/npg/Documents/bigrips_summary.sh which can be modified and used for future experiments. |
|